OpenStudy (anonymous):

Find the slope of the line that passes through the points (-2, 4) and (-5, -6). -2/7 (10/3) correct? -2/3

OpenStudy (anonymous):

y goes down by 10 for every 3 steps of x going down, so yes, 10/3 makes sense.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

slope = (y2 - y10 / (x2 - x1) slope = (-6-4) / (-5 - (-2) slope = -`10/(-5 + 2) slope = -10/-3 slope = 10/3 you are correct

OpenStudy (anonymous):

Find the slope of the line that passes through the points (-3, -1) and (0, 3). 3/4 0 (4/3) correct?

OpenStudy (anonymous):

Use the same logic as kellie showed you in the post above and you should be able to tell yourself i guess

OpenStudy (anonymous):

slope = (3 - (-1) / (0 - (-3) slope = (3 + 1) / (0 + 3) slope = 4/3 correct

OpenStudy (anonymous):

thanks. I really need help on this equation. Input in standard form the equation of the given line. The line that passes through (1, 1) and (3, 4). I know I'm wrong but my answer is y3-x2=1